5. Parexel Co. produces and sells cell phone batteries. Its production function is 𝒒 = πŸ‘πŸπŸπ‘²πŸŽ.πŸ‘π‘³πŸŽ.πŸ‘, while its cost function is 𝒄 = πŸπŸπŸŽπ‘² + πŸ—πŸŽπ‘³. Parexel can sell its batteries at 𝒑 = πŸ”πŸ‘.

a. Does this production function exhibit increasing, decreasing or constant returns to scale? Explain briefly.

[2]

b. Does this production function exhibit decreasing returns to capital? Explain briefly. [2]

c. Use implicit differentiation to find the marginal rate of technical substitution, 𝑴𝑹𝑻𝑺. [2]

d. Write down an expression for the isoquant if 𝒒 = πŸπŸŽπŸ’πŸ–. [2]

e. Write down an expression for the labour elasticity of production, 𝒆𝑳. [2]

f. Which levels of 𝑳 and 𝑲 satisfy the the first-order and second-order conditions for the maximisation of

Parexel’s profit? [8]

g. Write down the levels of profit, revenue, cost and output at the profit maximising levels of 𝑲 and 𝑳. [2]


1
Expert's answer
2021-06-24T12:36:10-0400

a

"q_0=321K^{0.3}L^{0.3}"

multiply each input by m

"q=321(m)^{0.6}K^{0.3}L^{0.3}\\m^{0.6}q_0"

i.e output increases by less than m, hence decreasing return to scale


b.


"MP_k=(321)(0.3)L^{0.3}K^{-0.7}\\\frac{\delta MP_k}{\delta k}=(321)(0.3)(-0.7)L^{0.3}K^{-1.7}<0"


This implies MPk is decreasing

hence Diminishing marginal return


c.

"MRTS=\frac{\frac{\delta q}{\delta L}}{\frac{\delta q}{\delta K}}"


"=\frac{(321)(0.3)K^{0.3}L^{-0.7}}{(321)(0.3)K^{-0.7}L^{0.3}}"


"MRTS=\frac{K}{L}"


d.

"q=2048\\\implies2048=321K^{0.3}L^{0.3}\\\implies K^{0.3}L^{0.3}\\\implies K^{0.3}L^{0.3}=6.4"Β Isoquant


e.

labor elasticity

"\in_L=(\frac{l}{q})(\frac{dq}{dl})"

"=\frac{l}{(321)K^{0.3}L^{0.3}}((321)(+0.3)K^{0.3}L^{-0.7})"


"\in=0.3"


f.

Profit will be mximized when

MPk=210

"(321)(0.3)K^{-0.7}L^{0.3}=210...................................(1)"

amd MPl=90

"(321)(0.3)K^{0.3}L^{-0.7}=-90..................................(2)"

byΒ "\frac{(1)}{(2)}"

"\frac{L}{K}=\frac{21}{9}=\frac{7}{3}"

"L=\frac{7}{3}K"

from cost function

"2048=321(K^{0.3})(\frac{7}{3})^{0.3}K^{0.3}\\\implies K^{0.6}=8.2\sqrt2"


"K=33.52\\L=78.22"


g.

"cost (C)=210K+90L\\=210(33.52)+90(78.23)= 1079"

"Revenue=P\times Q= (63)(2048)

=129024\\profit=TR-TC\implies129024-1079=14945"

From the question q is given as 2048 therefore,

"\\output=2048"